a vat contain 100L of a 40% acid solution. how much acid must be added to increase the acidity to 60% ?
Answer by checkley71(8403)

(X+100).6=100(.4)+X OR .6X+60=40+X OR 20=.4X OR 50=X
PROOF 150 TIMES .6=40+50 OR 90=90